Rob has twice as much as Sam. Then Sam gives Rob 1 quarter, 2 nickels, and 3 pennies. Rob then gives Sam 8 dimes. If they both end up with the same amount. How much did Rob originally have?
Answer by Edwin McCravy(20086)

Rob begins with R cents, Sam begins with S cents.

Rob has twice as much as Sam.
R = 2S

Then Sam gives Rob 1 quarter, 2 nickels, and 3 pennies.
(That's 38 cents)
Now Sam has S-38 cents, and Rob has R+38

Rob then gives Sam 8 dimes.
(That's 80 cents)
Now Sam has S-38+80, and Rob has R+38-80.
Simplifying,
Now Sam has S+42 cents and Rob has R-42  

...they both end up with the same amount.
 
(set them equal)
So  S+42 = R-42
       S = R-84
So we have this system of equations to solve:
R = 2S
S = R-84
Solve that system by substitution and get 
R = 168 cents or $1.68, S = 84 cents.
Rob originally had $1.68
Edwin